OpenStudy (anonymous):

Please help! To find the height of a pole, a surveyor moves 140 feet away from the base of the pole and then, with a transit 4 feet tall, measures the angle of elevation to the top of the pole to be 44 degrees. To the nearest foot, what is the height of the pole? I'm learning about vectors right now. Options: o 145 feet o 149 feet o 135 feet o 139 feet

OpenStudy (anonymous):

 145..Correct me if i am wrong in interpreting the question.

OpenStudy (anonymous):

I used tan. Tan(44) = x/140 x = 140(Tan 44) x = 135.196 The transit is 4, so we add that to 135. This makes 139 feet. Does this seem right to you?

OpenStudy (anonymous):

Ya absolutely right if the tan44 value you used is right. Sorry I do not have a calculator and so made approximations..

OpenStudy (anonymous):

Okay, thanks! :) Would you mind helping me out with one more problem? It's the last one I have to do.

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Miguel is driving his motorboat across a river. The speed of the boat in still water is 14 mi/h. The river flows directly south at 2 mi/h. If Miguel heads directly west, what are the boat's resultant speed and direction?

OpenStudy (anonymous):



OpenStudy (anonymous):

Do you know how to find a resultant vector? and its magnitude?

OpenStudy (anonymous):

Yes, I believe so.

OpenStudy (anonymous):

Then you have to just find the magnitude of two vectors here. One vector is 14 west and the other is 2 south. You have already found the direction of the resultant vector as southwest. Now find magnitude. r = sqrt(p^2 + q^2 + 2pqcosx) where r is the resultant vector' smagnitude, p and q are the known magnitudes and x is the angle between p and q. In this case, x=90. So the expression becomes r = sqrt( 14^2 + 2^2) Find this value..i think it will be 10sqrt(2)

OpenStudy (anonymous):

I did that and got c^2 = 196 + 4. c ^2 = the square root of 200 This comes out to be 14.14. That's the speed. Then I used Tan. Tan x degrees = 2/14 Inverse tan(2/14) = x x = 8.13 That's the direction. What do you think?

OpenStudy (anonymous):

Yea man..you know everything. Just build your confidence levels. GOOD WORK!

OpenStudy (anonymous):

Haha thanks! :)

OpenStudy (anonymous):

Have a nice day!